Integrate using the fractional fracture development (point b). For a and c, first divide by the division angle.

a.) ∫3x2+2x+2/x-3 dx

b.) ∫3/x2-1 dx

c.) ∫2x2+x/x+2 dx
1
Expert's answer
2020-01-20T10:15:41-0500

Solution.

a. We write the integral in the form


"\intop \frac {3x^2+2x+2}{x-3}dx=\intop \frac {3x(x-3)+11x+2}{x-3}dx="

"=\intop \frac {3x(x-3)+11(x-3)+35}{x-3}dx= \intop (3x+11+\frac{35} {x-3})dx"

Using the tabular values of the antiderivative we get

"\intop \frac {3x^2+2x+2}{x-3}dx= \frac{3x^2}{2}+11x+35 ln |x-3| +C"

where C is constant.

b. Represent the fraction 3/x2-1 as the sum of two fractions


"\frac {3} {x^2-1}=\frac{A}{x-1} + \frac {B}{x+1}= \frac{Ax+A+Bx-B}{x^2-1}"

Comparing the coefficients near the powers of x we get the system of equations


"\begin{cases}
 A+B=0 \\
 A-B=3
\end{cases} \implies \begin{cases}
 A=\frac{3}{2} \\
 B=-\frac {3}{2}
\end{cases}"

We write the integral in the form

"\intop \frac {3}{x^2-1}dx=\frac {3}{2} \intop (\frac{1}{x-1}-\frac{1}{x+1})dx"

Using the tabular values of the antiderivative we get


"\intop \frac {3}{x^2-1}dx=\frac{3}{2} ln|x-1|-\frac{3}{2} ln|x+1|+C=\frac{3}{2} ln \frac {|x-1|}{|x+1|}+C"

where C is constant.

c. We write the integral in the form


"\intop \frac {2x^2+x}{x+2}dx=\intop \frac {2x(x+2)-3x}{x+2}dx="

"\intop \frac {2x(x+2)-3(x+2) +6}{x+2}dx=\intop (2x-3+\frac {6}{x+2})dx"

Using the tabular values of the antiderivative we get


"\intop \frac {2x^2+x}{x+2}dx=x^2-3x+6ln|x+2|+C"

where C is constant.

Answer. a.


"\intop \frac {3x^2+2x+2}{x-3}dx= \frac{3x^2}{2}+11x+35 ln |x-3| +C"

where C is constant.

b.


"\intop \frac {3}{x^2-1}dx=\frac{3}{2} ln \frac {|x-1|}{|x+1|}+C"

where C is constant.

c.


"\intop \frac {2x^2+x}{x+2}dx=x^2-3x+6ln|x+2|+C"

where C is constant.